OpenStudy (anonymous):

Two automobiles leave from the same point in Chicago at the same time and travel along straight highways that differ by 65o. If their speeds are 62 mi/hr and 74 mi/hr, respectively, determine how far apart the cars are after 29 minutes. Round answer to two decimal places.

OpenStudy (anonymous):

65o?

OpenStudy (anonymous):

here are the options:40.76 miles B. 25.48 miles C. 45.86 miles D. 30.57 miles E. 35.66 miles

OpenStudy (anonymous):

35.66 miles is the answer. Note that in 29 minutes the cars travel 62 x (29/60) = 29.6666667 miles and 74 x (29/60) = 35.76666667 miles then using the Law of Cosines we have c^2 = 29.67^2 + 35.77^2 - 2(29.67)(35.77)cos(65) and taking the square root we have c = 35.66 miles